OpenStudy (anonymous):

Find the exact solution for (√(5) - 1)/(x) = (√(5))/(2) Then, find the approximate solution.

OpenStudy (jack1):

(√(5) - 1)/(x) = (√(5))/(2) √(5) - 1 = (√(5))/(2) x √(5) - 1 = (√(5))/(2) x √(5) = (√(5))/(2) x + 1 √(5) - (√(5))/(2) = x + 1 (√(5))/(2) = x + 1 x = (√(5))/(2) - 1

OpenStudy (anonymous):

Thank you so much :)